Ann is four years younger than Maria. In three years, ann is two thirds Maria's age. How old are they now?
Answer by macston(5194)

A=Ann's age; M=Maria's age=A+4yrs
A+3yrs=2/3(M+3yrs) Substitute for M
A+3yrs=2/3(A+4yrs+3yrs) Multiply each side by 3.
3A+9yrs=2A+14yrs Subtract 2A+9 from each side
A=5yrs ANSWER 1: Ann is 5 years old.
M=A+4yrs=5yrs+4yrs=9yrs Maria is 9 years old.
CHECK:
A+3yrs=2/3(M+3yrs)
5yrs+3yrs=2/3(9yrs+3yrs)
8yrs=2/3(12yrs)
8yrs=8yrs